OpenStudy (anonymous):

I'll tell you a secret if you help me? :))))) If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation.8x - 2y = 4 4x + 5y = -18 Answer options: 4x + 5(-4x - 2) = -18 4x + 5(4x - 2) = -18 4(-4x - 2) + 5y = -18 4(4x - 2) + 5y = -18

OpenStudy (venomblast):

bet

OpenStudy (venomblast):

leave the y by itself

OpenStudy (venomblast):

well you can solve for y. iof you divide -2 on both side, y should equal -2

OpenStudy (venomblast):

now plug -2 for y

OpenStudy (anonymous):

I have that :)

OpenStudy (anonymous):

help me!!!!!!!!!!!:(((((((((((((

OpenStudy (venomblast):

ok ok lol

OpenStudy (venomblast):

if you plug in for y. x should equal 2

OpenStudy (anonymous):

The answer I got was C

OpenStudy (anonymous):

no worries dear :) @iheartfood by the way, I <3 food too!!!

OpenStudy (venomblast):

so 4(2) + 5(-4(2) - 2) = -18 8+5(-8-2)=-18 8+5(-10)=-18 8-50=-18 -42=-18 FALSE

OpenStudy (venomblast):

now all you have to do is plug it in for each. dont worry i will help you

OpenStudy (anonymous):

awwwww you are sooo sweeeeet!!!!! :))))))))) @Venomblast

OpenStudy (venomblast):

hehe thanks :)

OpenStudy (anonymous):

Can you help me finish the problem with how you were taught so I can refer back to it and learn? :))

OpenStudy (venomblast):

yea. but i already told you. all you have to do is find what x and y equal and then plug it in.

OpenStudy (anonymous):

Then which would be the answer is what I'm asking...my error

OpenStudy (anonymous):

:(

OpenStudy (venomblast):

ok the 1st one wont work. i tried the second and still doesnt work

OpenStudy (anonymous):

I told you the answer I got was C

OpenStudy (venomblast):

you did? i didnt get it i think. imaybe i just miss it

OpenStudy (anonymous):

It's fine love

OpenStudy (venomblast):

for C 4(-4x - 2) + 5y = -18 4((-4(-2))-2)+5(-2)+=-18 4(8-2)-10=-18 4(6)-10=-18 24-10=-18 FALSE

OpenStudy (anonymous):

OMG I GET IT NOW!!!!!! THANK YOU!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! i CORRECTED MY ERRORS

OpenStudy (anonymous):

YOURE THE BEST!!!!! SERIOUSLY!!!!!!!!!

OpenStudy (venomblast):

what was wrong? did you do a typo?

OpenStudy (venomblast):

lol what did i do?

OpenStudy (venomblast):

haha im lost :)

OpenStudy (anonymous):

YOU HELPED ME:) I checked my work and found miniscule errors

OpenStudy (venomblast):

that great :D

OpenStudy (anonymous):

god you're so sweet! I could hug you xoxoxoxox

OpenStudy (venomblast):

if you were in NYC lol i can

OpenStudy (venomblast):

I love to help. I am also going to become a math tutor. paid.

OpenStudy (anonymous):

I wish! It being as famous as it is and all :)

OpenStudy (venomblast):

im really happy you got your solution :).

OpenStudy (anonymous):

you'd make a terrific one, seriously... you're so understanding and thorough

OpenStudy (anonymous):

you'd make for a great partner

OpenStudy (venomblast):

awww that was sweet what you said. i feel like taking a snapshot. you being my reference will convince people i am great at math

OpenStudy (venomblast):

any more questions ?:)

OpenStudy (anonymous):

awwwwww! I'd love to tell them great things about your SUPERIOR math skills!

OpenStudy (venomblast):

hehe :) im flattered :)

OpenStudy (anonymous):

I have MANY more unfortunatly, but i don't want to bother you...

OpenStudy (venomblast):

well im here to help. there no such thing as bothering if your on a site like this. i like helping people even if its for free

OpenStudy (anonymous):

you're such a kind being. I like to help people on yahoo answers and elsewhere, but when it comes to math, I have very LITTLE to offer, so I really do appreciate your efforts

OpenStudy (venomblast):

hehe it cool. i can help :)

OpenStudy (anonymous):

The sad part is I'm in a rush... I have so many questions and very little time

OpenStudy (anonymous):

I'm very very very scared

OpenStudy (venomblast):

why? math test?

OpenStudy (anonymous):

And I've literally been upset and crying all day. My eyes have been blood shot red.

OpenStudy (venomblast):

oh :( if this get personal to you, just send me a personal email :) just for your privacy. what happen by the way. :o

OpenStudy (anonymous):

you can email me if you want some time :) i have a lot to offer. part of what makes me upset about getting on here, is that people know me for my lack of understanding in math, and that's the only judgement they can make about me, when i'm very intelligent in other aspects

OpenStudy (venomblast):

do people judge you at this site?

OpenStudy (anonymous):

I've gotten very negative feedback a few times

OpenStudy (anonymous):

Cyber bullying... you know how that goes

OpenStudy (venomblast):

in this site?!?!?!

OpenStudy (anonymous):

in fact, today I heard that I was "retarded" on here.. it made me feel OH SO WARM INSIDE. not

OpenStudy (venomblast):

really.. i thought this was an peaceful site.

OpenStudy (anonymous):

another time this one lady gave me the hardest time when all I was doing was learning. i could easily say "f you" to learning math, but I try to at least get my answers checked and look for explanations if i'm incorrect

OpenStudy (venomblast):

yea dont bother with people that has no patience. not worth it, i can make you learn math. i love to help

OpenStudy (anonymous):

you're really a great person. I'm not even kidding you.

OpenStudy (anonymous):

I don't want to keep you up though... what time is it in NYC?

OpenStudy (venomblast):

12:54 am :D it alright

OpenStudy (anonymous):

You're an hour ahead of me

OpenStudy (venomblast):

you live in Central time right?

OpenStudy (anonymous):

But considering NY is such a lively place, i'm sure you're used to being up late, and yes

OpenStudy (venomblast):

yea you got that right. the city that never sleep :)

OpenStudy (anonymous):

If you help me fast, I'll literally do anything. Time is ticking and i would never want to use someone and make them feel unwanted, but I want and need you right now.

OpenStudy (venomblast):

ok :) ask away

OpenStudy (anonymous):

i don't even have time for explanations as to why the answer is right. I know that sounds horrible. but I am f******

OpenStudy (venomblast):

lol take a deep breath and relax lol. you already type fast. that a good start. so tell me the question.

OpenStudy (venomblast):

from this question or a new question?

OpenStudy (anonymous):

hahahahahahaha thank you!!!! ;) i fkn' owe you

OpenStudy (venomblast):

nah it free of charge :P

OpenStudy (anonymous):

I still owe you with something. i'll keep in touch with you and let you in on my mysterious ways

OpenStudy (venomblast):

hmmm. im curious now lol

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

1. Solve the following system of equations.2x + 3y - z = 13x + y + 2z = 12x + 2y - 3z = -5 (3, 1, 2) (-3, 1, 2) (3, -1, 2) (3, 1, -2)

OpenStudy (venomblast):

that one whole equation?

OpenStudy (venomblast):

let me guess, the 1st column is x the second is y and the last is z?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I think you break it down into two, so you separate it

OpenStudy (venomblast):

lol i got this

OpenStudy (anonymous):

into three rather***

OpenStudy (anonymous):

I have faith in YOU;)

OpenStudy (venomblast):

is that a 2x or .2x?

OpenStudy (anonymous):

huh?

OpenStudy (venomblast):

on the beginning of the question

OpenStudy (anonymous):

No period in the beginning, my error 2x + 3y - z = 13x + y + 2z = 12x + 2y - 3z = -5

OpenStudy (venomblast):

what math course are you taking?

OpenStudy (anonymous):

Algebra

OpenStudy (anonymous):

I guess you can't help me....

OpenStudy (venomblast):

the answer is A

OpenStudy (venomblast):

lol i took calculus and im not even on that class yet

OpenStudy (venomblast):

hello? the answer is A